OpenStudy (anonymous):

David made this model of a pyramid. Which option best describes the length of the unknown side of David’s pyramid? Answer A)less than 4 cm B)greater than 24 cm C)between 4 cm and 20 cm D)between 20 cm and 24 cm

OpenStudy (anonymous):

would you attach the pyramid?

OpenStudy (precal):



OpenStudy (precal):

use a^2+b^2=c^2 a^2+8^2=12^2 solve for a